# SiDiaC: Sinhala Diachronic Corpus

### v1

@@ 4.3 Evaluation of the Corpus
SiDiaC consists of 58,027 word tokens that were filtered using regex, retaining only Sinhala and Latin characters, and subsequently tokenised by whitespace. The corpus contains 833 words in Latin script, which accounts for just 1.42% of the entire dataset. Also, the complete dataset comprises 22,837 unique word tokens in Sinhala script, which accounts for 39.36% unique word coverage of all words. This total word token count, while not in the range of millions, such as the COHA corpus for English, comfortably passes the 53,000 token count of FarPaHC for Faroese, which is in the same language resource category as Sinhala according to Ranathunga and de Silva (2022).
@@
Additionally, we conducted an analysis to identify the stopwords in the corpus by examining word tokens at the century level. Following the method described by Wijeratne and de Silva (2020) for their contemporary Sinhala corpus, we used a combination of word frequency analysis and manual vetting to identify appropriate stopwords from the corpora. To achieve this, we first counted the frequency of each unique word for each century and then converted these frequencies into z-scores.
@@
The stopword that ranked highest across six of the seven centuries was ‘ [Uncaptioned image] ’(was, became), while ‘ [Uncaptioned image] ’(or, either) ranked highest in the fifth century. Additionally, other frequently occurring stopwords include [Uncaptioned image] (that, if), ‘ [Uncaptioned image] ’(that[named]), ‘ [Uncaptioned image] ’(that, those), ‘ [Uncaptioned image] ’(this, these), ‘ [Uncaptioned image] ’(or, interrogative particle), ‘ [Uncaptioned image] ’(while), ‘ [Uncaptioned image] ’(by), and ‘ [Uncaptioned image] ’(be [exists]). However, it was observed that certain words, such as ‘ [Uncaptioned image] ’(honorific for a revered person) and ‘ [Uncaptioned image] ’(Quality) were also receiving high z-scores, potentially due to the corpus’s strong connection to Buddhist literature. The methodology for identifying the top 48 stopwords in the entire corpus and their depiction is presented in Appendix C.
@@ 5 Conclusion
The complete corpus was thoroughly analysed, highlighting the powerful OCR performance of Document AI beyond simple text extraction. This was followed by a detailed evaluation of the dataset based on the metadata of all the books. Additionally, a comprehensive analysis was conducted at the word token level to ensure the identification of important findings within the corpus.
In future studies, this dataset can be further enhanced by using more literary works and improving its quality through better post-processing techniques. It provides capabilities to study language evolution, including lexical semantic change, the appearance of new words (neologisms), and corpus-based lexicography, among other applications.
@@
Literature Identification: While we recognised the Department of National Archives [^21] https://websnew.lithium.lk/archives/ of Sri Lanka also as a credible source, data acquisition was conducted only from the National Library of Sri Lanka due to permission constraints.
Data Filtration: Out of the 221 scanned copies acquired, we were able to identify the written dates or periods for only 59 of them. The written dates of the books were annotated based on the lifespans of well-known authors, while the majority of the remainder were annotated relying heavily on the work by  Sannasgala (2015), which represents an overreliance on a single source.
Post-Processing after OCR: Under this process, corrections were initiated to address identified formatting issues, while possible identification errors at the word or character level were not addressed.
Code-Mixed Data: It was noted that the corpus contains code mixing of Pali, Sanskrit, and English languages, but the removal of text in these languages from the corpus has not been done.
Commentary Books: The identified books primarily named with the term ‘ [Uncaptioned image] ’ (meaning commentaries) will include two written dates for the original and the commentary. However, these instances were anchored to the well-known original version without removing the commentary.
Appendix A Post-Processing Extracted Text
@@
[…]
It was notable to see errors in spacing in poems, especially when the last word or letter of a word is separated by multiple spaces that were not correctly detected by Document AI, as shown in Table 3. Although correcting these errors may not hold significant semantic importance, accurately replicating the structure of the poems is crucial for preserving the original form of the books in case any downstream task using our dataset requires it. But for any task that does not require the original structure and would work on the lexical or semantic properties of the writing, the Document AI output is adequate.
There were also multi-column texts, particularly evident in poetry books such as ‘Kavya Wajrayudhaya - Palamu Kotasa’. These texts contained errors where two columns were treated as a single column, resulting in entire horizontal lines being extracted without properly traversing each column, as shown in Table 4. To address these errors, corrections were made by mimicking the original structure of the books to preserve their intended format. The corrections provided in the table demonstrate that OCR outputs can be restructured faithfully only through column-aware preprocessing (for example, layout analysis, region detection, or image segmentation).
The misplaced words and phrases appeared multiple times due to varying spacing and text positioning styles used in different books. Unlike simple character substitutions, these errors dislocate words or phrases from their expected syntactic and semantic slots. For example, text fragments are merged across lines, while key markers or paragraph boundaries vanish. This was particularly noticeable at the beginning of the first paragraph in texts with drop caps. Similarly, in poetry, irregular line breaks caused by extra spaces before the last letter or word were among these issues. A misplaced suffix or dislocated phrase could invert rhetorical emphasis or obscure reference chains. Importantly, these errors also impact computational parsing, where algorithms expecting consistent lineation and phrase boundaries will misinterpret discourse structure. These issues are illustrated in Table 5. Another one of the most common issues addressed during post-processing was the indentation errors in paragraphs and lines. These errors were often found in the first line of paragraphs, with certain poems being misaligned, and page headings being centred incorrectly. A few examples of these issues are presented in Table 5.
Document AI also captured meta information on the physical book such as page numbers. During the text extraction, these were removed from the books because they do not add any value to the presented corpus. These numbers were typically placed at the top or bottom, and they were centred or right-aligned in different literary works.
@@
Anagathawanshaya: Methe Budu Siritha
Watadhdhara Medhanandha Himi;
Siri Parakumabahu;
Wilgammula Sangaraja Himi
Fiction Religious 1934 1325 - 1333 0.9992
@@
Gadaladeni Sannayai Prasidhdha wu
Balawathare Purana Wyakyanaya
Hikkaduwe Sri Sumangala Himi Non-Fiction Language 1877 1827 - 1911 0.9970
@@
Dharma Pradheepikawa hewath
Mahabodhiwansha Parikathawa
Unknown Non-Fiction Religious 1906 1187 - 1225 0.9682
@@
Nikaya Sangrahaya hewath Shasanawatharaya
Unknown
Non-Fiction Religious 1922 1390 0.9754
Nidhahase Manthraya
S Mahinda Himi
Non-Fiction Poetry 1938 1901 - 1938 0.8997
Pansiya Panas Jathaka Potha
Unknown
Fiction Religious 1881 1303 - 1333 0.9987
Parawi Sandheshaya
Unknown
Fiction Poetry 1873 1430 - 1440 0.9902
Parani Gama
Galpatha Kemanandha Himi
Non-Fiction History 1944 1944 0.9846
Budhdha Sikka hewath Kudu Sika
Unknown
Non-Fiction History 1898 1270 - 1293 0.9699
Mage Malli
G. H Perera
Non-Fiction Poetry 1938 1886 - 1938 0.8659
Mahawansa Teeka
Hikkaduwe Sri Sumangala Himi
Non-Fiction History 1895 1827 - 1895 0.9978
Muwadew da Wiwaranaya
Munidhasa Kumarathunga
Non-Fiction Religious 1949 1887 - 1944 0.8710
Moggalalayanawyakaranan
Moggallana Himi
Non-Fiction Language 1890 1070 - 1232 0.9051
Moggallana Panchika Pradeepaya
Unknown
Non-Fiction Language 1896 1070 - 1232 0.9918
Liyanora Nadagama
Unknown
Fiction Poetry 1936 1852 - 1927 0.9935
Wibath Maldhama
Kirama Dhammarama Himi
Non-Fiction Language 1906 1821 0.9986
Waidya Chinthamani Baishadhya Sangrahawa
Unknown
Non-Fiction Medical 1909 1706 - 1739 0.9965
Wyakarana Wiwarana hewath
Sinhala Bashawe Wyakaranaya
Munidhasa Kumarathunga
Non-Fiction Language 1937 1887 - 1937 0.9029
Sadhdharma Rathnawaliya -Prathama Bagaya
Dharmasena Himi
Non-Fiction Religious 1930 1220 - 1293 0.9962
Sanna sahitha Abhisambodhi Alankaraya
Waliwita Saranankara Sangaraja Himi
Non-Fiction Religious 1897 1698 - 1778 0.9989
Sanna sahitha Salalihini Sandheshaya
Unknown
Fiction Religious 1859 1450 0.9909
Sanskrutha Shabdhamalawa
hewath Sanskrutha Nama Waranagilla
Rathmalane Dharmaloka Himi
Non-Fiction Language 1876 1828 - 1887 0.9671
Sarartha Sangrahawa: Prathama Bhagaya
Srimadh Budhdhadhasa Rajathuma
Non-Fiction Medical 1904 398 - 426 0.9997
Sithiyam sahitha Mahiyangana Warnanawa
Unknown
Fiction Poetry 1898 1878 0.9989
Sithiyam sahitha Sadhdharmalankaraya
Unknown
Non-Fiction Religious 1954 1398 - 1410 0.9810
Sithiyam sahitha Siyabas Maldhama
Kirama Dhammanandha Himi
Fiction Poetry 1894 1820 0.9256
Sithiyam sahitha Sinhala Mahawanshaya
D.H.S Abhayarathna
Non-Fiction History 1922 1874 0.9549
Sinhala Wyakaranaya enam Sidath Sangarawa
Hikkaduwe Sri Sumangala Himi
Non-Fiction Language 1884 1827 - 1911 0.9886
Hansa Sandheshaya
C.E. Godakumbure
Fiction Poetry 1953 1457 - 1465 0.8553
Hithopadhesha Sannaya
Waligama Sri Sumangala Himi
Non-Fiction
- 1884 1825 - 1905 0.9871
@@ C.E. Godakumbure
Appendix B Utilized Literary Works
The metadata includes the title in Sinhala, the romanised title, the author’s name, the romanised author’s name, the genre, the issue date, the writing date, and the OCR confidence level. The complete metadata for each literary work used in this compilation of the dataset can be found in Table 6 with the titles and authors’ names presented in romanised Sinhala. This diachronic spread ensures coverage of Sinhala evolution across medieval, pre-modern, and modern stages. Religious texts dominate, reflecting both preservation biases and the centrality of Buddhism in Sinhala literary culture.
@@ Appendix B Utilized Literary Works
Appendix C Analysis of Stop Words in SiDiaC
@@
The continued inclusion of particles and suffixes shows continuity of core grammatical function words. Their presence across the 5th–20th centuries suggests remarkable diachronic stability in Sinhala morphosyntactic scaffolding. However, the table also reveals anomalies: certain lexical items not traditionally classified as stopwords appear in the stopword list. This distortion is likely due to corpus composition (religious texts with Buddhist themes dominate some centuries, inflating the relative frequency of doctrinal terms). Another observation is the persistence of poetic suffixes in older centuries, gradually tapering in modern texts. This diachronic shift may point to a movement away from poetry and towards prose. The presence/absence patterns also reveal data sparsity in some centuries, as marked by gaps where certain stopwords do not appear due to limited surviving texts. To wit, it is as interesting (mayhap more) to note what is missing in the 20th century, given that it seems to count a majority of the candidate words among its stopwords. Note how the archaic forms of honorifics and some direct references to Buddhism have dropped out of the list. In earlier centuries, the preponderance of monastic authorship and the dominance of canonical or exegetical works meant that words indexing reverence and religious entities were unavoidable high-frequency items. Their disappearance, or at least their reduced prominence, in the most recent century may be reflecting how the subjects covered in the text have shifted from esoteric religious communication to comparatively more secular discourse.

### v2

@@ 4.3 Evaluation of the Corpus
SiDiaC consists of 58,027 word tokens that were filtered using regex, retaining only Sinhala and Latin characters, and subsequently tokenised by whitespace. The corpus contains 833 words in Latin script, which accounts for just 1.42% of the entire dataset. Also, the complete dataset comprises 22,837 unique word tokens in Sinhala script, which accounts for 39.36% unique word coverage of all words. This total word token count, while not in the range of millions, such as the COHA corpus for English, comfortably passes the 53,000 token count of FarPaHC for Faroese, which is in the same language resource category as Sinhala according to ranathunga-de-silva-2022-languages.
@@
Additionally, we conducted an analysis to identify the stopwords in the corpus by examining word tokens at the century level. Following the method described by wijeratne2020sinhala for their contemporary Sinhala corpus, we used a combination of word frequency analysis and manual vetting to identify appropriate stopwords from the corpora. To achieve this, we first counted the frequency of each unique word for each century and then converted these frequencies into z-scores.
$z_{w,c}=\frac{f_{w,c}-\mu_{c}}{\sigma_{c}}$ (1)
where, $z_{w,c}$ is the z-score of the word $w$ in the century $c$. The term $f_{w,c}$ refers to the frequency of the word $w$ during that century, $\mu_{c}$ denotes the mean frequency of all words in century $c$, and $\sigma_{c}$ indicates the standard deviation of the frequencies of all words in that century.
@@
The stopword that ranked highest across six of the seven centuries was ‘ [Uncaptioned image] ’(was, became), while ‘ [Uncaptioned image] ’(or, either) ranked highest in the fifth century. Additionally, other frequently occurring stopwords include [Uncaptioned image] (that, if), ‘ [Uncaptioned image] ’(that[named]), ‘ [Uncaptioned image] ’(that, those), ‘ [Uncaptioned image] ’(this, these), ‘ [Uncaptioned image] ’(or, interrogative particle), ‘ [Uncaptioned image] ’(while), ‘ [Uncaptioned image] ’(by), and ‘ [Uncaptioned image] ’(be [exists]). However, it was observed that certain words, such as ‘ [Uncaptioned image] ’(honorific for a revered person) and ‘ [Uncaptioned image] ’(Quality) were also receiving high z-scores, potentially due to the corpus’s strong connection to Buddhist literature. The methodology for identifying the top 48 stopwords in the entire corpus and their depiction is presented in Appendix D.
5 Future Work
SiDiaC, being the first resource of its kind, paves the way for diachronic linguistic studies of the Sinhala language. We encourage researchers to explore areas such as lexical semantic change, tracking neologisms, grammatical change, historical language modelling, and corpus-based lexicography. Additionally, since the data is annotated by genre, it also allows for synchronic studies focusing on domain-based differences.
The corpus, as mentioned earlier, has comfortably surpassed the 53,000-token count of FarPaHC for Faroese, which falls into the same language resource category according to ranathunga-de-silva-2022-languages. However, the token count of SiDiaC, currently at 58,027, could certainly be increased by adding more literary works and additional pages from the existing 46 books. This enhancement would support more accurate research studies.
It is important to note that the OCR post-processing conducted in this study focuses only on formatting. However, as discussed in Appendix E, there are clear issues present at the word and character levels that need to be meticulously addressed. Furthermore, as mentioned earlier, the corpus is code-mixed with Pali, Sanskrit, and English. This highlights the need for a processing step to identify and remove irrelevant text, ensuring that the corpus is entirely focused on Sinhala.
We also identified that books called ‘ [Uncaptioned image] ’ (meaning commentaries) may include two dates: one for the original (quoted) text and another for the commentary. However, this study did not consider this phenomenon, and such instances were attributed to the well-known original version. Therefore, in future studies, it would be beneficial to identify the two dates and include the text sections originating from the different time periods at their correct positions in the corpus.
The SiDiaC corpus did not undergo any lexical annotations during this study. Typically, the most recognised method for creating diachronic corpora involves parsing the entire corpus using POS tagging. However, this approach was not feasible with Sinhala POS taggers due to their limited performance. In future research, it would be highly beneficial to have the entire corpus parsed manually by Sinhala linguists who understand the evolution of language structure in Sinhala.
@@ 6 Conclusion
The complete corpus was thoroughly analysed, highlighting the powerful OCR performance of Document AI beyond simple text extraction. This was followed by a detailed evaluation of the dataset based on the metadata of all the books. Additionally, a comprehensive analysis was conducted at the word token level to ensure the identification of important findings within the corpus. Finally, we discussed potential future studies and approaches that could enhance the dataset, as well as the research opportunities that this corpus provides.
@@
Literature Identification:
While we recognised the Department of National Archives [^21] https://websnew.lithium.lk/archives/ of Sri Lanka also as a credible source, data acquisition was conducted only from the National Library of Sri Lanka due to permission constraints.
Data Filtration:
Out of the 221 scanned copies acquired, we were able to identify the written dates or periods for only 59 of them. The written dates of the books were annotated based on the lifespans of well-known authors, while the majority of the remainder were annotated relying heavily on the work by  Sannasgala_2009, which represents an overreliance on a single source.
Post-Processing after OCR:
Under this process, while corrections were initiated to address identified formatting issues, possible identification errors at the word or character level discussed in Appendix E were not addressed.
Code-Mixed Data:
It was noted that the corpus contains code mixing of Pali, Sanskrit, and English languages, but the removal of text in these languages from the corpus has not been done.
Commentary Books:
The identified books primarily named with the term ‘ [Uncaptioned image] ’ (meaning commentaries) will include two written dates for the original and the commentary. However, these instances were anchored to the well-known original version without removing the commentary.
Lexical Annotation:
Unlike the LatinlSE, IcePAHC, COHA, and Google N-gram corpora, which have undergone lexical annotations specifically for POS tagging, we were unable to conduct similar annotations due to the unavailability of Sinhala POS taggers de2025survey.
Acknowledgments
The creation of the SiDiaC corpus was made possible through the valuable contributions of several individuals. We extend our sincere gratitude to Padma Bandaranayake, Director of the National Library & Documentation Centre, for her assistance with data acquisition. We also acknowledge Uthpala Nimanthi and Charani Palangasinghe for their efforts in the post-processing of the data, and the expertise of Nalaka Jayasena, a Sinhala Linguist, which was important in the identification of the book by Sannasgala_2009. Finally, we would like to thank Jayath de Silva, Savin Madapatha, and Thushan Bawantha for their dedicated work on the written date annotation.
Appendix A Utilised Literary Works
The metadata includes the title in Sinhala, the romanised title, the author’s name, the romanised author’s name, the genre, the issue date, the writing date, and the OCR confidence level.
The complete metadata for each literary work used in this compilation of the dataset can be found in Table 3 with the titles and authors’ names presented in romanised Sinhala. This diachronic spread ensures coverage of Sinhala evolution across medieval, pre-modern, and modern stages. Religious texts dominate, reflecting both preservation biases and the centrality of Buddhism in Sinhala literary culture.
Appendix B Comparison of Document AI & Surya
The thorough quantitative analysis conducted by jayatilleke2025zero indicates that Surya outperforms Document AI when evaluated on a synthetically created Sinhala dataset. However, during our text extraction process, we found that Document AI actually surpasses Surya. We believe this discrepancy stems from the synthetic data used in their study, which does not accurately reflect the challenges presented by real scanned documents.
Table 4 highlights three examples that clearly demonstrate why Document AI is the superior OCR engine. The errors indicated in red boxes for both systems demonstrate that Document AI excels in character identification, particularly with diacritics and similar-looking letters. Additionally, Document AI appears to be the only system effectively implementing morpheme segmentation, which is crucial for maintaining consistent word forms over time. Lastly, Document AI’s text modernisation feature provides another significant advantage, making it the ideal choice for integration into the OCR pipeline used in this study.
Appendix C Post-Processing Extracted Text
@@
Anagathawanshaya: Methe Budu Siritha Watadhdhara Medhanandha Himi; Siri Parakumabahu; Wilgammula Sangaraja Himi Fiction Religious 1934 1325 - 1333 0.9992
@@
Gadaladeni Sannayai Prasidhdha wu Balawathare Purana Wyakyanaya Hikkaduwe Sri Sumangala Himi Non-Fiction Language 1877 1827 - 1911 0.9970
@@
Dharma Pradheepikawa hewath Mahabodhiwansha Parikathawa Unknown Non-Fiction Religious 1906 1187 - 1225 0.9682
@@
Nikaya Sangrahaya hewath Shasanawatharaya Unknown Non-Fiction Religious 1922 1390 0.9754
Nidhahase Manthraya S Mahinda Himi Non-Fiction Poetry 1938 1901 - 1938 0.8997
Pansiya Panas Jathaka Potha Unknown Fiction Religious 1881 1303 - 1333 0.9987
Parawi Sandheshaya Unknown Fiction Poetry 1873 1430 - 1440 0.9902
Parani Gama Galpatha Kemanandha Himi Non-Fiction History 1944 1944 0.9846
Budhdha Sikka hewath Kudu Sika Unknown Non-Fiction History 1898 1270 - 1293 0.9699
Mage Malli G. H Perera Non-Fiction Poetry 1938 1886 - 1938 0.8659
Mahawansa Teeka Hikkaduwe Sri Sumangala Himi Non-Fiction History 1895 1827 - 1895 0.9978
Muwadew da Wiwaranaya Munidhasa Kumarathunga Non-Fiction Religious 1949 1887 - 1944 0.8710
Moggalalayanawyakaranan Moggallana Himi Non-Fiction Language 1890 1070 - 1232 0.9051
Moggallana Panchika Pradeepaya Unknown Non-Fiction Language 1896 1070 - 1232 0.9918
Liyanora Nadagama Unknown Fiction Poetry 1936 1852 - 1927 0.9935
Wibath Maldhama Kirama Dhammarama Himi Non-Fiction Language 1906 1821 0.9986
Waidya Chinthamani Baishadhya Sangrahawa Unknown Non-Fiction Medical 1909 1706 - 1739 0.9965
Wyakarana Wiwarana hewath Sinhala Bashawe Wyakaranaya Munidhasa Kumarathunga Non-Fiction Language 1937 1887 - 1937 0.9029
Sadhdharma Rathnawaliya -Prathama Bagaya Dharmasena Himi Non-Fiction Religious 1930 1220 - 1293 0.9962
Sanna sahitha Abhisambodhi Alankaraya Waliwita Saranankara Sangaraja Himi Non-Fiction Religious 1897 1698 - 1778 0.9989
Sanna sahitha Salalihini Sandheshaya Unknown Fiction Religious 1859 1450 0.9909
Sanskrutha Shabdhamalawa hewath Sanskrutha Nama Waranagilla Rathmalane Dharmaloka Himi Non-Fiction Language 1876 1828 - 1887 0.9671
Sarartha Sangrahawa: Prathama Bhagaya Srimadh Budhdhadhasa Rajathuma Non-Fiction Medical 1904 398 - 426 0.9997
Sithiyam sahitha Mahiyangana Warnanawa Unknown Fiction Poetry 1898 1878 0.9989
Sithiyam sahitha Sadhdharmalankaraya Unknown Non-Fiction Religious 1954 1398 - 1410 0.9810
Sithiyam sahitha Siyabas Maldhama Kirama Dhammanandha Himi Fiction Poetry 1894 1820 0.9256
Sithiyam sahitha Sinhala Mahawanshaya D.H.S Abhayarathna Non-Fiction History 1922 1874 0.9549
Sinhala Wyakaranaya enam Sidath Sangarawa Hikkaduwe Sri Sumangala Himi Non-Fiction Language 1884 1827 - 1911 0.9886
Hansa Sandheshaya C.E. Godakumbure Fiction Poetry 1953 1457 - 1465 0.8553
Hithopadhesha Sannaya Waligama Sri Sumangala Himi Non-Fiction - 1884 1825 - 1905 0.9871
@@
Table 4: Examples of sentences along with their corresponding text extractions from Surya and Document AI for comparison. Note that the characters and phrases highlighted in red boxes contain errors. $\dagger$ This character is known as ‘ [Uncaptioned image] ’, a punctuation mark that indicates the end of a text or section in historical Sinhala.
[…]
@@
It was notable to see errors in spacing in poems, especially when the last word or letter of a word is separated by multiple spaces that were not correctly detected by Document AI, as shown in Table 5. Although correcting these errors may not hold significant semantic importance, accurately replicating the structure of the poems is crucial for preserving the original form of the books in case any downstream task using our dataset requires it. But for any task that does not require the original structure and would work on the lexical or semantic properties of the writing, the Document AI output is adequate. There were also multi-column texts, particularly evident in poetry books such as ‘Kavya Wajrayudhaya - Palamu Kotasa’. These texts contained errors where two columns were treated as a single column, resulting in entire horizontal lines being extracted without properly traversing each column, as shown in Table 6. To address these errors, corrections were made by mimicking the original structure of the books to preserve their intended format.
The corrections provided in the table demonstrate that OCR outputs can be restructured faithfully only through column-aware preprocessing (for example, layout analysis, region detection, or image segmentation).
The misplaced words and phrases appeared mul-
-tiple times due to varying spacing and text positioning styles used in different books. Unlike simple character substitutions, these errors dislocate words or phrases from their expected syntactic and semantic slots. For example, text fragments are merged across lines, while key markers or paragraph boundaries vanish. This was particularly noticeable at the beginning of the first paragraph in texts with drop caps. Similarly, in poetry, irregular line breaks caused by extra spaces before the last letter or word were among these issues. A misplaced suffix or dislocated phrase could invert rhetorical emphasis or obscure reference chains. Importantly, these errors also impact computational parsing, where algorithms expecting consistent lineation and phrase boundaries will misinterpret discourse structure. Another one of the most common issues addressed during post-processing was the indentation errors in paragraphs and lines. These errors were often found in the first line of paragraphs, with certain poems being misaligned, and page headings being centred incorrectly. A few examples of these issues are presented in Table 7.
Document AI also captured meta information on the physical book, such as page numbers. During the text extraction, these were removed from the books because they do not add any value to the presented corpus. These numbers were typically placed at the top or bottom, and they were centred or right-aligned in different literary works.
Appendix D Analysis of Stop Words in SiDiaC
@@
The continued inclusion of particles and suffixes shows continuity of core grammatical function words. Their presence across the 5th–20th centuries suggests remarkable diachronic stability in Sinhala morphosyntactic scaffolding. However, the table also reveals anomalies: certain lexical items not traditionally classified as stopwords appear in the stopword list. This distortion is likely due to corpus composition (religious texts with Buddhist themes dominate some centuries, inflating the relative frequency of doctrinal terms). Another observation is the persistence of poetic suffixes in older centuries, gradually tapering in modern texts.
This diachronic shift may point to a movement away from poetry and towards prose. The presence/absence patterns also reveal data sparsity in some centuries, as marked by gaps where certain stopwords do not appear due to limited surviving texts. To wit, it is as interesting (mayhap more) to note what is missing in the 20th century, given that it seems to count a majority of the candidate words among its stopwords. Note how the archaic forms of honorifics and some direct references to Buddhism have dropped out of the list.
In earlier centuries, the preponderance of monastic authorship and the dominance of canonical or exegetical works meant that words indexing reverence and religious entities were unavoidable high-frequency items. Their disappearance, or at least their reduced prominence, in the most recent century may be reflecting how the subjects covered in the text have shifted from esoteric religious communication to comparatively more secular discourse. Equally important is the fact that the persistence of other grammatical function words across all centuries stands in stark contrast to this attrition of religiously marked lexemes. This points to a kind of lexical stratification: the unmarked syntactic scaffolding of Sinhala remains stable over time (only being replaced by synonyms when they do), while the culturally bound vocabulary tied to ritual, doctrine, or honorific practice is more vulnerable to historical change.
Appendix E Word & Character Level Errors
During the post-processing of extracted text from scanned PDF files, we carefully conducted formatting level corrections as mentioned in section 3.6. However, these adjustments did not resolve all the necessary corrections at the word and character levels.
Table 9: Examples of word deformation caused by character-level identification errors, including *incorrect identifications of diacritics/letters and **a complete deletion of a character at the marker. $\dagger$ Note that this is not an error, but rather a text modernisation step.
It became clear that character identification issues persisted throughout the documents. As illustrated in Table 9, diacritics in Sinhala posed significant challenges. Some characters were completely unrecognised, resulting in character deletions. Additionally, some identified diacritics were incorrect substitutions for different diacritics. There were also instances where a diacritic was erroneously added even when no corresponding character existed, indicative of character insertion errors. In simpler terms, most of the errors in this category are related to spelling issues.